# Linear-Cost Vecchia Approximation of Multivariate Normal Probabilities

### v4

@@ 1 Introduction
Multivariate normal (MVN) probabilities, which are integrals over MVN densities, must be evaluated in a variety of settings, most notably those involving Gaussian processes (GPs), including GP regression with inequality constraints (Da Veiga & Marrel, 2012), joint threshold-exceedance probabilities for GPs (Bolin & Lindgren, 2015), probit GP models for binary responses (Durante, 2019), skew-normal models (Zhang & El-Shaarawi, 2010), and censored scale-mixtures of normals for modeling tail dependence (Huser et al., 2017). However, MVN probabilities are analytically intractable, and direct application of numerical integration rules often leads to highly inaccurate estimates (Genz, 1992).
@@
For evaluating high-dimensional MVN probabilities with spatial covariance matrices, Genton et al. (2018) and Cao et al. (2021) used hierarchical-matrix approximations to reduce the complexity of pre-processing and MC sampling for SOV by a factor of $n^{1/2}$. Bolin & Lindgren (2015) re-parameterized the SOV method with the inverse Cholesky factor to take advantage of the sparse precision matrix under Gaussian Markov random fields. However, they found this approach to often be computationally inefficient, because the sparsity was not well preserved during Cholesky factorization. Other recent approaches for spatial MVN probabilities have used versions of the Vecchia approximation (Vecchia, 1988), which decomposes a multivariate density into a product of univariate conditional densities with small conditioning sets. Saha et al. (2022) applied a Vecchia-type approximation to reduce the complexity of computing the Cholesky factor in the preprocessing stage, but its per-sample complexity remains as $\mathcal{O}(n^{2})$. Nascimento & Shaby (2022) decomposed the $n$-dimensional MVN probability into a product of lower-dimensional MVN probability ratios by assuming conditional independence in the MVN cumulative distribution function (CDF); while this is based on a similar principle as the Vecchia approximation, it appears that the screening effect (i.e., far-away information contributes little after conditioning on nearby points) (Stein, 2002) is weaker for MVN CDFs than for densities, resulting in substantially increased approximation error for CDFs as will be shown empirically in this paper.
Our proposed methods are also based on the Vecchia approximation. A Vecchia approximation of a MVN distribution results in a valid MVN distribution that has a covariance matrix with a sparse inverse Cholesky factor (Datta et al., 2016; Katzfuss & Guinness, 2021). Vecchia approximations can provide asymptotically exact inference in near-linear time for GPs with Matérn kernels subject to boundary effects and smoothness conditions (Schäfer et al., 2021; Kang et al., 2024) — such kernels and their relation to sparse precision matrices were also studied in Lindgren et al. (2011). Numerically, Vecchia approximations have been shown to be highly accurate even with small conditioning-set sizes in many settings, including spatial statistics and high-dimensional and multi-output GPs (e.g., Datta et al., 2016; Katzfuss et al., 2022; Schäfer et al., 2021; Cao et al., 2023). Vecchia approximations can be applied to any covariance matrix, and we expect them to be most accurate if the covariance can be viewed as a stationary kernel with relatively small noise-to-signal ratio defined over some (potentially unknown) input space (Kang & Katzfuss, 2023).
@@ 3.1 Conditional mean and variance under Vecchia
where $\eta_{i}=E[x_{i}|\mathbf{x}_{c(i)}]$ and $l_{i}^{2}=\mbox{Var}[x_{i}|\mathbf{x}_{c(i)}]$ can be computed straightforwardly based on an $(m+1)\times(m+1)$ submatrix of $\bm{\Sigma}$. Each conditioning set is typically chosen as the indices of the $m$ nearest previously ordered neighbors. For covariances $\bm{\Sigma}_{i,j}=\mathcal{K}(\mathbf{s}_{i},\mathbf{s}_{j})$ based on isotropic kernels, Euclidean distance between inputs $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ can be used to define nearest neighbors; for other covariance matrices, the correlation distance $(1-|\rho_{ij}|)^{1/2}$ with $\rho_{ij}=\bm{\Sigma}_{ij}/(\bm{\Sigma}_{ii}\bm{\Sigma}_{jj})^{1/2}$ implicitly selects neighbors in a suitable transformed input space (Kang & Katzfuss, 2023). The Vecchia approximation is based on the screening effect (Stein, 2002), which describes the phenomenon that the marginal information gain from far-away responses is often minimal after conditioning on nearby responses. Schäfer et al. (2021) showed that, for each $i=1,\ldots,n$, the nonzero entries in the $i$th column of the inverse Cholesky factor $\mathbf{V}=\mathbf{L}^{-\top}$ can be obtained as:
@@
where $\mathbf{A}$ is a sparse matrix $\mathbf{A}^{\top}=(diag(\mathbf{l})-\mathbf{V})\cdot diag(\mathbf{l})^{-1}$, $\mathbf{V}$ is the (sparse) inverse Cholesky factor of $\bm{\Sigma}$, $\bm{\Sigma}=(\mathbf{V}\mathbf{V}^{\top})^{-1}$, and $\mathbf{l}$ is the inverse of the diagonal entries of $\mathbf{V}$.
Proposition 2 is similar to existing results (e.g., Rue & Held, 2005, Thm. 2.3). Its proof is included in the appendix for completeness. The definition of $\mathbf{l}$ in Proposition 2 coincides with that above, namely the univariate conditional standard deviations. Notice that our definitions of $\mathbf{A}$ and $\mathbf{l}$ share similarities with Bolin & Lindgren (2015), which re-parameterized the SOV method with the inverse Cholesky factor, but Bolin & Lindgren (2015) computed $\mathbf{V}$ from the sparse precision matrix, hence leading to higher complexity and less sparsity in $\mathbf{V}$. The conditional expectations and variances computed as proposed in Proposition 2 result in biased MVN probabilities. However, the bias is shown in Proposition 3 to decrease exponentially with the number of non-zero entries in $\mathbf{V}$ under certain regularity conditions, covering Matérn kernels with smoothness $\nu$ such that $\nu+d/2$ is an integer, where $d$ is the dimension of the domain over which Matérn kernels are defined (Schäfer et al., 2021).
@@ 4 Solving For Exponential Tilting Parameters
Based on these expressions, we propose to minimize $g$ for VMET using a gradient-based optimizer, for which $\nabla g=\mathbf{H}\cdot\nabla\psi$ can be computed efficiently in $\mathcal{O}(n)$ time, and which does not require the more expensive $\mathbf{H}^{-1}\cdot\nabla\psi$. Specifically, due to the sparsity of $\mathbf{A}$ with fewer than $nm$ non-zero coefficients, $\nabla\psi$ in (18) and matrix-vector products involving $\mathbf{H}$ can be computed in $\mathcal{O}(nm)$ time. In our implementation, we use the limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm (Liu & Nocedal, 1989) as our optimizer, which creates a low-rank approximation of the Hessian of $g$, maintaining an overall $\mathcal{O}(n)$ complexity. Another advantage of VMET over MET is that the Hessian of $\psi$ with respect to $\mathbf{x}$, $\frac{\partial^{2}\psi}{\partial\mathbf{x}^{2}}$, can be represented using the sparse matrix $\mathbf{A}$, and so $\operatorname*{arg\,max}_{\mathbf{x}}\psi(\mathbf{x};\bm{\gamma})$ can be solved efficiently with second-order optimization algorithms. Finding $\max_{\mathbf{x}}\psi(\mathbf{x};\bm{\gamma})$, which is the logarithm of the maximum PDF ratio (i.e., $\max\log h$ in (9)), accurately for a given $\bm{\gamma}$ is crucial to the accept-reject sampling of TMVN distributions introduced in Section 2.
@@ 5.1 Variable Reordering Based on Vecchia
Variable reordering can be reduced to linear complexity under a special case of the Vecchia approximation with $c(i)=\{1,\ldots,m\}$ for $i>m$, which results in the fully independent conditional (FIC) approach (Snelson & Ghahramani, 2007; Katzfuss & Guinness, 2021), a popular GP approximation in machine learning. Under the FIC assumption, the variable ordering (after selecting the first $m$ variables) can be obtained in a single iteration, hence achieving an overall $\mathcal{O}(n)$ complexity. For the first $m$ variables, we can simply run the loop in Algorithm LABEL:alg:vecc_reorder for $m$ iterations. We refer to this $\mathcal{O}(n)$ ordering as the FIC-based variable reordering to distinguish it from the $\mathcal{O}(n^{2})$ Vecchia-based variable reordering in Algorithm LABEL:alg:vecc_reorder. We provide a numerical comparison of the univariate reordering, the Vecchia-based variable reordering, and the FIC-based variable reordering in Section 6.1.
@@ 6 Numerical comparison
the SOV method from Genz (1992); Genz & Bretz (2009)
@@
the Vecchia approximation of MVN CDFs from Nascimento & Shaby (2022)
The original implementation of SOV in the R package mvtnorm (Genz & Bretz, 2009) only allows input MVN dimensions $n$ up to 1,0001{,}000; for n>1,000n>1{,}000 we used the SOV implementation in tlrmvnmvt (Cao et al., 2022). The implementations used for MET, TLR, and VCDF were the TruncatedNormal (Botev, 2017) R package, the tlrmvnmvt (Cao et al., 2022) R package, and the R package from https://github.com/Recca2012/CDFApprox, respectively. tlrmvnmvt, TruncatedNormal, and our implemented VMET all utilized C/C++ source code through Rcpp (Eddelbuettel & François, 2011), while VCDF is based on a pre-compiled binary package. Computation times were measured on a standard scientific workstation.
@@ 6.1 Variable reordering comparison
In this section, we show that the variable reordering heuristic proposed for SOV (Gibson et al., 1994) can also improve the results for MET and VMET. We demonstrate that our proposed Vecchia-based variable reordering (Algorithm LABEL:alg:vecc_reorder) can achieve the same level of improvement in reducing the MC variance as the univariate reordering (Algorithm LABEL:alg:univar_reorder). SOV and TLR are not considered in this section, because their benefits from variable reordering have been documented in the literature (e.g., Genz & Bretz, 2009; Cao et al., 2022).
@@ 6.3 Comparison for high-dimensional MVN probabilities
For Scenarios 1 and 2, we considered VMET with $m=70$ as the benchmark, as VMET arguably has lower MC variance than SOV and TLR for tail MVN probabilities, and the bias caused by the Vecchia approximation was likely insignificant compared with the MC variance. For Scenario 3, we used SOV as benchmark, because for centered MVN probabilities, MET and SOV amount to the same integrand, hence SOV is expected to have the same accuracy as MET. Compared to SOV and TLR, VMET significantly reduced the MC variance, hence improving estimation accuracy for tail MVN probabilities; for centered MVN probabilities, VMET also achieved a more favorable trade-off between estimation error and computation cost. Figure LABEL:fig:high_dim_exp_boxplot in the Appendix provides more detailed information on the three methods’ probability estimates. Furthermore, TLR, the state-of-the-art in scalability, may suffer from unpredictable numerical singularity, whereas VMET is numerically stable, because unlike low-rank matrix approximations, the Vecchia approximation has a theoretical guarantee for positive definiteness (see Prop. 1 in Katzfuss & Guinness, 2021).
@@ 7.1 Parameter estimation
It is common practice in many environmental applications to substitute censored data with a constant value, such as half the level of detection (LOD), the LOD divided by the square root of 2, or simply zero (Croghan & Egeghy, 2003). These LOD-based methods are also broadly used as a benchmark for more sophisticated Markov-chain-Monte-Carlo-based methods (De Oliveira, 2005; Ordoñez et al., 2018), which are often not scalable to large $n$. We show that in the presence of censored data, the censored MVN model in (21) can achieve more accurate parameter estimation than LOD-based methods. We generated a spatial field $\mathbf{x}$ over 6,4006{,}400 locations on a $80\times 80$ grid in the unit square based on a GP with mean zero and a Matérn covariance with variance, range, smoothness, and nugget of $(1.0,0.1,1.5,0.03)$ in the same setup as Scenario 1 in Table 1. All $x_{i}$ values below $b_{i}=0$ were censored to obtain the data $\mathbf{z}$. Because the censoring thresholds were zero, all LOD-based methods discussed in (Croghan & Egeghy, 2003) amounted to the same data augmentation.
@@ 7.3 Analysis of groundwater contamination
The responses and detection thresholds after transformation were modeled by a partially censored MVN distribution with zero mean and a Matérn-1.5 kernel. We used one range parameter for longitude and latitude and another one for the temporal coordinate. Four parameters $(\sigma^{2},\beta_{1},\beta_{2},\tau^{2})$, namely variance, spatial range, temporal range, and nugget, were estimated by maximizing the likelihood (21) as described in Section 7.1. Notice that we used the chordal distance for our study since the spatial locations are all within the United States, where the chordal distance closely resembles the great-circle distance. For optimization, we set $m=50$, MC sample size to $10^{4}$, and used the robust Nelder and Mead Algorithm (Nelder & Mead, 1965) native to R, which took approximately five hours to reach convergence; the parameters were estimated to be $(13.28,0.04,12.46,0.07)$, indicating very strong temporal correlation.

### v5

@@ 1 Introduction
Multivariate normal (MVN) probabilities, which are integrals over MVN densities, must be evaluated in a variety of settings, most notably those involving Gaussian processes (GPs), including GP regression with inequality constraints (Da Veiga and Marrel, 2012), joint threshold-exceedance probabilities for GPs (Bolin and Lindgren, 2015), probit GP models for binary responses (Durante, 2019), skew-normal models (Zhang and El-Shaarawi, 2010), and censored scale-mixtures of normals for modeling tail dependence (Huser et al., 2017). However, MVN probabilities are analytically intractable, and direct application of numerical integration rules often leads to highly inaccurate estimates (Genz, 1992).
@@
For evaluating high-dimensional MVN probabilities with spatial covariance matrices, Genton et al. (2018) and Cao et al. (2021) used hierarchical-matrix approximations to reduce the complexity of pre-processing and MC sampling for SOV by a factor of $n^{1/2}$. Bolin and Lindgren (2015) re-parameterized the SOV method with the inverse Cholesky factor to take advantage of the sparse precision matrix under Gaussian Markov random fields. However, they found this approach to often be computationally inefficient, because the sparsity was not well preserved during Cholesky factorization. Other recent approaches for spatial MVN probabilities have used versions of the Vecchia approximation (Vecchia, 1988), which decomposes a multivariate density into a product of univariate conditional densities with small conditioning sets. Saha et al. (2022) applied a Vecchia-type approximation to reduce the complexity of computing the Cholesky factor in the preprocessing stage, but its per-sample complexity remains as $\mathcal{O}(n^{2})$. Nascimento and Shaby (2022) decomposed the $n$-dimensional MVN probability into a product of lower-dimensional MVN probability ratios by assuming conditional independence in the MVN cumulative distribution function (CDF); while this is based on a similar principle as the Vecchia approximation, it appears that the screening effect (i.e., far-away information contributes little after conditioning on nearby points) (Stein, 2002) is weaker for MVN CDFs than for densities, resulting in substantially increased approximation error for CDFs as will be shown empirically in this paper.
Our proposed methods are also based on the Vecchia approximation. A Vecchia approximation of a MVN distribution results in a valid MVN distribution that has a covariance matrix with a sparse inverse Cholesky factor (Datta et al., 2016; Katzfuss and Guinness, 2021). Vecchia approximations can provide asymptotically exact inference in near-linear time for GPs with Matérn kernels subject to boundary effects and smoothness conditions (Schäfer et al., 2021; Kang et al., 2024) — such kernels and their relation to sparse precision matrices were also studied in Lindgren et al. (2011). Numerically, Vecchia approximations have been shown to be highly accurate even with small conditioning-set sizes in many settings, including spatial statistics and high-dimensional and multi-output GPs (e.g., Datta et al., 2016; Katzfuss et al., 2022; Schäfer et al., 2021; Cao et al., 2023). Vecchia approximations can be applied to any covariance matrix, and we expect them to be most accurate if the covariance can be viewed as a stationary kernel with relatively small noise-to-signal ratio defined over some (potentially unknown) input space (Kang and Katzfuss, 2023).
@@ 3.1 Conditional mean and variance under Vecchia
where $\eta_{i}=E[x_{i}|\mathbf{x}_{c(i)}]$ and $l_{i}^{2}=\mbox{Var}[x_{i}|\mathbf{x}_{c(i)}]$ can be computed straightforwardly based on an $(m+1)\times(m+1)$ submatrix of $\bm{\Sigma}$. Each conditioning set is typically chosen as the indices of the $m$ nearest previously ordered neighbors. For covariances $\bm{\Sigma}_{i,j}=\mathcal{K}(\mathbf{s}_{i},\mathbf{s}_{j})$ based on isotropic kernels, Euclidean distance between inputs $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ can be used to define nearest neighbors; for other covariance matrices, the correlation distance $(1-|\rho_{ij}|)^{1/2}$ with $\rho_{ij}=\bm{\Sigma}_{ij}/(\bm{\Sigma}_{ii}\bm{\Sigma}_{jj})^{1/2}$ implicitly selects neighbors in a suitable transformed input space (Kang and Katzfuss, 2023). The Vecchia approximation is based on the screening effect (Stein, 2002), which describes the phenomenon that the marginal information gain from far-away responses is often minimal after conditioning on nearby responses. Schäfer et al. (2021) showed that, for each $i=1,\ldots,n$, the nonzero entries in the $i$th column of the inverse Cholesky factor $\mathbf{V}=\mathbf{L}^{-\top}$ can be obtained as:
@@
where $\mathbf{A}$ is a sparse matrix $\mathbf{A}^{\top}=(diag(\mathbf{l})^{-1}-\mathbf{V})\cdot diag(\mathbf{l})$, $\mathbf{V}$ is the (sparse) inverse Cholesky factor of $\bm{\Sigma}$, $\bm{\Sigma}=(\mathbf{V}\mathbf{V}^{\top})^{-1}$, and $\mathbf{l}$ is the inverse of the diagonal entries of $\mathbf{V}$.
Proposition 2 is similar to existing results (e.g., Rue and Held, 2005, Thm. 2.3). Its proof is included in the appendix for completeness. The definition of $\mathbf{l}$ in Proposition 2 coincides with that above, namely the univariate conditional standard deviations. Notice that our definitions of $\mathbf{A}$ and $\mathbf{l}$ share similarities with Bolin and Lindgren (2015), which re-parameterized the SOV method with the inverse Cholesky factor, but Bolin and Lindgren (2015) computed $\mathbf{V}$ from the sparse precision matrix, hence leading to higher complexity and less sparsity in $\mathbf{V}$. The conditional expectations and variances computed as proposed in Proposition 2 result in biased MVN probabilities. However, the bias is shown in Proposition 3 to decrease exponentially with the number of non-zero entries in $\mathbf{V}$ under certain regularity conditions, covering Matérn kernels with smoothness $\nu$ such that $\nu+d/2$ is an integer, where $d$ is the dimension of the domain over which Matérn kernels are defined (Schäfer et al., 2021).
@@ 4 Solving For Exponential Tilting Parameters
Based on these expressions, we propose to minimize $g$ for VMET using a gradient-based optimizer, for which $\nabla g=\mathbf{H}\cdot\nabla\psi$ can be computed efficiently in $\mathcal{O}(n)$ time, and which does not require the more expensive $\mathbf{H}^{-1}\cdot\nabla\psi$. Specifically, due to the sparsity of $\mathbf{A}$ with fewer than $nm$ non-zero coefficients, $\nabla\psi$ in (18) and matrix-vector products involving $\mathbf{H}$ can be computed in $\mathcal{O}(nm)$ time. In our implementation, we use the limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm (Liu and Nocedal, 1989) as our optimizer, which creates a low-rank approximation of the Hessian of $g$, maintaining an overall $\mathcal{O}(n)$ complexity. Another advantage of VMET over MET is that the Hessian of $\psi$ with respect to $\mathbf{x}$, $\frac{\partial^{2}\psi}{\partial\mathbf{x}^{2}}$, can be represented using the sparse matrix $\mathbf{A}$, and so $\operatorname*{arg\,max}_{\mathbf{x}}\psi(\mathbf{x};\bm{\gamma})$ can be solved efficiently with second-order optimization algorithms. Finding $\max_{\mathbf{x}}\psi(\mathbf{x};\bm{\gamma})$, which is the logarithm of the maximum PDF ratio (i.e., $\max\log h$ in (9)), accurately for a given $\bm{\gamma}$ is crucial to the accept-reject sampling of TMVN distributions introduced in Section 2.
@@ 5.1 Variable Reordering Based on Vecchia
Variable reordering can be reduced to linear complexity under a special case of the Vecchia approximation with $c(i)=\{1,\ldots,m\}$ for $i>m$, which results in the fully independent conditional (FIC) approach (Snelson and Ghahramani, 2007; Katzfuss and Guinness, 2021), a popular GP approximation in machine learning. Under the FIC assumption, the variable ordering (after selecting the first $m$ variables) can be obtained in a single iteration, hence achieving an overall $\mathcal{O}(n)$ complexity. For the first $m$ variables, we can simply run the loop in Algorithm LABEL:alg:vecc_reorder for $m$ iterations. We refer to this $\mathcal{O}(n)$ ordering as the FIC-based variable reordering to distinguish it from the $\mathcal{O}(n^{2})$ Vecchia-based variable reordering in Algorithm LABEL:alg:vecc_reorder. We provide a numerical comparison of the univariate reordering, the Vecchia-based variable reordering, and the FIC-based variable reordering in Section 6.1.
@@ 6 Numerical comparison
the SOV method from Genz (1992); Genz and Bretz (2009)
@@
the Vecchia approximation of MVN CDFs from Nascimento and Shaby (2022)
The original implementation of SOV in the R package mvtnorm (Genz and Bretz, 2009) only allows input MVN dimensions $n$ up to 1,0001{,}000; for n>1,000n>1{,}000 we used the SOV implementation in tlrmvnmvt (Cao et al., 2022). The implementations used for MET, TLR, and VCDF were the TruncatedNormal (Botev, 2017) R package, the tlrmvnmvt (Cao et al., 2022) R package, and the R package from https://github.com/Recca2012/CDFApprox, respectively. tlrmvnmvt, TruncatedNormal, and our implemented VMET all utilized C/C++ source code through Rcpp (Eddelbuettel and François, 2011), while VCDF is based on a pre-compiled binary package. Computation times were measured on a standard scientific workstation.
@@ 6.1 Variable reordering comparison
In this section, we show that the variable reordering heuristic proposed for SOV (Gibson et al., 1994) can also improve the results for MET and VMET. We demonstrate that our proposed Vecchia-based variable reordering (Algorithm LABEL:alg:vecc_reorder) can achieve the same level of improvement in reducing the MC variance as the univariate reordering (Algorithm LABEL:alg:univar_reorder). SOV and TLR are not considered in this section, because their benefits from variable reordering have been documented in the literature (e.g., Genz and Bretz, 2009; Cao et al., 2022).
@@ 6.3 Comparison for high-dimensional MVN probabilities
For Scenarios 1 and 2, we considered VMET with $m=70$ as the benchmark, as VMET arguably has lower MC variance than SOV and TLR for tail MVN probabilities, and the bias caused by the Vecchia approximation was likely insignificant compared with the MC variance. For Scenario 3, we used SOV as benchmark, because for centered MVN probabilities, MET and SOV amount to the same integrand, hence SOV is expected to have the same accuracy as MET. Compared to SOV and TLR, VMET significantly reduced the MC variance, hence improving estimation accuracy for tail MVN probabilities; for centered MVN probabilities, VMET also achieved a more favorable trade-off between estimation error and computation cost. Figure LABEL:fig:high_dim_exp_boxplot in the Appendix provides more detailed information on the three methods’ probability estimates. Furthermore, TLR, the state-of-the-art in scalability, may suffer from unpredictable numerical singularity, whereas VMET is numerically stable, because unlike low-rank matrix approximations, the Vecchia approximation has a theoretical guarantee for positive definiteness (see Prop. 1 in Katzfuss and Guinness, 2021).
@@ 7.1 Parameter estimation
It is common practice in many environmental applications to substitute censored data with a constant value, such as half the level of detection (LOD), the LOD divided by the square root of 2, or simply zero (Croghan and Egeghy, 2003). These LOD-based methods are also broadly used as a benchmark for more sophisticated Markov-chain-Monte-Carlo-based methods (De Oliveira, 2005; Ordoñez et al., 2018), which are often not scalable to large $n$. We show that in the presence of censored data, the censored MVN model in (21) can achieve more accurate parameter estimation than LOD-based methods. We generated a spatial field $\mathbf{x}$ over 6,4006{,}400 locations on a $80\times 80$ grid in the unit square based on a GP with mean zero and a Matérn covariance with variance, range, smoothness, and nugget of $(1.0,0.1,1.5,0.03)$ in the same setup as Scenario 1 in Table 1. All $x_{i}$ values below $b_{i}=0$ were censored to obtain the data $\mathbf{z}$. Because the censoring thresholds were zero, all LOD-based methods discussed in (Croghan and Egeghy, 2003) amounted to the same data augmentation.
@@ 7.3 Analysis of groundwater contamination
The responses and detection thresholds after transformation were modeled by a partially censored MVN distribution with zero mean and a Matérn-1.5 kernel. We used one range parameter for longitude and latitude and another one for the temporal coordinate. Four parameters $(\sigma^{2},\beta_{1},\beta_{2},\tau^{2})$, namely variance, spatial range, temporal range, and nugget, were estimated by maximizing the likelihood (21) as described in Section 7.1. Notice that we used the chordal distance for our study since the spatial locations are all within the United States, where the chordal distance closely resembles the great-circle distance. For optimization, we set $m=50$, MC sample size to $10^{4}$, and used the robust Nelder and Mead Algorithm (Nelder and Mead, 1965) native to R, which took approximately five hours to reach convergence; the parameters were estimated to be $(13.28,0.04,12.46,0.07)$, indicating very strong temporal correlation.